what is the difference between molar mass and ‘1mole of atoms’? in youtube video i found the formula 1mole of atoms=molar mass/avogadro no.
please explain

Grade:Select Grade

1 Answers

Aarti Gupta
8 years ago
Mole Concept – If we take the smallest sample of a substance then it may contains a very large number of atoms or molecules.These atoms and molecules are so small in size that it is not possible to count them individually.So in order to overcome such difficulties,the conceprt of mole was introduced.According to which the no.of particles of the substance is related to mass of the substance.Mole may be defined as--
The amount of substance that contains as many specified elementary particles as the number of atoms in 12g of carbon-12 isotope.These elementary particles may be atoms,molecules,ions,electrons etc.Thus mole is a unit for counting such particles.It stands for 6.023 * 1023 particles irrespesctive of their nature.This no. is called as Avogadro’s no. or Avogadro’s constant.
Thus- 1 mole = 6.023 * 1023 particles.For ex- 1 mole of oxygen atoms = 6.023 * 1023 oxygen atoms.
1mole ofoxygen molecules = 6.023 * 1023 oxygen molecules
Molar mass– The mass of one mole of particles is called as molar mass.The number of atoms or molecules contained in molar mass is equal to 6.023 * 1023.For ex-
One mole of Na atoms means 6.023 * 1023 Na atoms weighing 23g which is the GAM(gram atomic mass) of Na.So we can write as -
Molar mass of Na = 23g/mol